OpenStudy (kayders1997):

Someone help find the integral of y/e^2y from 0 to 1

OpenStudy (kayders1997):

@Luigi0210

OpenStudy (kayders1997):

@zepdrix

OpenStudy (kayders1997):

@agent0smith

OpenStudy (kayders1997):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

have you tried integration by parts?

OpenStudy (kayders1997):

I'm not good at it :(

jimthompson5910 (jim_thompson5910):

Do you agree that \[\LARGE \frac{y}{e^{2y}} = y*e^{-2y}\] ??

OpenStudy (kayders1997):

Yes

OpenStudy (kayders1997):

Y would be e^2y would be easier to integrate

jimthompson5910 (jim_thompson5910):

I wanted to get it into u*dv form

OpenStudy (kayders1997):

So u is y? Right

jimthompson5910 (jim_thompson5910):

it might be easier to make u = e^(-2y) and dv = y I'm not 100% sure which path is easier

OpenStudy (kayders1997):

Oh okay :)

jimthompson5910 (jim_thompson5910):

let's try it out though u = e^(-2y) du = -2e^(-2y)dy dv = y v = (1/2)*y^2 agree with everything so far?

OpenStudy (kayders1997):

Yes

OpenStudy (kayders1997):

Uv-integral vdu

jimthompson5910 (jim_thompson5910):

this you mean? \[\LARGE \int u*dv = u*v - \int v*du\]

OpenStudy (kayders1997):

Correct

jimthompson5910 (jim_thompson5910):

So we have \[\large \int u*dv = u*v - \int v*du\] \[\large \int e^{-2y}*y = e^{-2y}*\left(\frac{1}{2}\right)*y^2 - \int \left(\frac{1}{2}\right)*y^2*(-2e^{-2y})dy\]

jimthompson5910 (jim_thompson5910):

hmm the problem looks like it's getting more complicated, let me try another way

OpenStudy (kayders1997):

Okay

OpenStudy (kayders1997):

Take your time I got a lot of time

jimthompson5910 (jim_thompson5910):

You were right in making u = y. That's much simpler u = y du = dy dv = e^(-2y)dy v = (-1/2)*e^(-2y) \[\Large \int u*dv = u*v - \int v*du\] \[\Large \int y*e^{-2y}dy = y*\left(-\frac{1}{2}*e^{-2y}\right) - \int \frac{-1}{2}*e^{-2y}dy\] \[\Large \int y*e^{-2y}dy = -\frac{1}{2}*y*e^{-2y} + \frac{1}{2}\int e^{-2y}dy\] Do you see how to finish up?

OpenStudy (kayders1997):

Is it -y2(e^-2y)-1/4e^-2y?

OpenStudy (kayders1997):

-y/2

jimthompson5910 (jim_thompson5910):

I think you mean this right? \[\Large -\frac{y}{2}e^{-2y}-\frac{1}{4}e^{-2y}+C\]

OpenStudy (kayders1997):

Yeah except it's a definite integral

jimthompson5910 (jim_thompson5910):

ok so you'll let \[\Large g(y) = -\frac{y}{2}e^{-2y}-\frac{1}{4}e^{-2y}+C\]

jimthompson5910 (jim_thompson5910):

then you'll compute the value of `g(1) - g(0)`

jimthompson5910 (jim_thompson5910):

the constant C will cancel when computing `g(1) - g(0)`

OpenStudy (kayders1997):

Here give me a few minutes and I'll let you know what I get :)

OpenStudy (kayders1997):

Okay -3/4e^-2-1/4?

jimthompson5910 (jim_thompson5910):

Yes the answer is \[\Large \frac{1}{4}-\frac{3}{4}e^{-2}\]

jimthompson5910 (jim_thompson5910):

I'm getting +1/4 not -1/4

OpenStudy (kayders1997):

Hmmmm

jimthompson5910 (jim_thompson5910):

plug in y = 0 \[\Large g(y) = -\frac{y}{2}e^{-2y}-\frac{1}{4}e^{-2y}+C\] \[\Large g(0) = -\frac{0}{2}e^{-2*0}-\frac{1}{4}e^{-2*0}+C\] \[\Large g(0) = -\frac{1}{4}+C\] plug in y = 1 \[\Large g(y) = -\frac{y}{2}e^{-2y}-\frac{1}{4}e^{-2y}+C\] \[\Large g(1) = -\frac{1}{2}e^{-2*1}-\frac{1}{4}e^{-2*1}+C\] \[\Large g(1) = -\frac{1}{2}e^{-2}-\frac{1}{4}e^{-2}+C\] now subtract \[\Large g(1) - g(0) = \left(-\frac{1}{2}e^{-2}-\frac{1}{4}e^{-2}+C\right) - \left(-\frac{1}{4}+C\right)\] \[\Large g(1) - g(0) = -\frac{1}{2}e^{-2}-\frac{1}{4}e^{-2}+C +\frac{1}{4}-C\] \[\Large g(1) - g(0) = -\frac{3}{4}e^{-2}+\frac{1}{4}\] \[\Large g(1) - g(0) = \frac{1}{4}-\frac{3}{4}e^{-2}\]

jimthompson5910 (jim_thompson5910):

you might have forgotten about a negative somewhere

OpenStudy (kayders1997):

Oh yeah it's minus a negative oops!

jimthompson5910 (jim_thompson5910):

other than that mistake, you did good

OpenStudy (kayders1997):

Thank you, and thank you for the help :)